# LEO-Mini: An Efficient Multimodal Large Language Model using Conditional Token Reduction and Mixture of Multi-Modal Experts

### v1

@@ 3.3 Mixture of Multi-modal Experts (MMoE)
3.4 Training stages
The training of LEO-Mini consists of three stages, as shown in Tab. 1. Stage 1: Warming up for the visual projector. This stage pre-trains the visual projector while keeping the LLM and vision experts frozen. The LLM and vision experts are initialized from the base models, while the vision projector is randomly initialized. Stage 2: Supervised fine-tuning. In this stage, we fine-tune the entire model, including the LLM, vision experts, and visual projector. Stage 3: Supervised fine-tuning for reducing the number of visual tokens. In this stage, we introduce CoTR to the model and perform LoRA fine-tuning (MMoE-LLM) for efficiency. Specifically, CoTR and MMoE-LLM are randomly initialized and fine-tuned, while the LLM and vision experts remain frozen. To avoid the risk of routing collapse [53], we employ a balanced loss for MMoE with a hyperparameter $\lambda=0.05$, following previous works [53, 56].
@@ 3.4 Training stages
Model # of General OCR Knowledge
[…]
VoCo-LLaMA [63] 1 1323.3 58.8 53.7 57.0 - - 65.4 81.4 -
LLaMA-VID [33] 2 - - - 55.5 - 49.0 67.7 83.1 -
PruMerge [52] 32 1350.3 60.9 - - - 56.0 72.0 76.3 -
MQT-LLaVA [20] 64 1464.3 63.5 - 60.0 51.5 - 67.0 83.6 34.4
Token-Packer [31] 64 - 64.1 - 61.1 50.7 - - 86.3 -
LLaVA-Mini [68] 64 1476.8 67.5 60.2 61.8 58.5 59.1 69.7 85.3 -
LEO-Mini-Vicuna-7B 64 1542.6 67.8 73.2 64.0 51.4 70.1 73.3 90.0 34.1
$\Delta$uparrow 65.8 $\uparrow$ 0.3 $\uparrow$ 13.0 $\uparrow$ 2.2 $\downarrow$ 7.1 $\uparrow$ 11.0 $\uparrow$ 1.3 $\uparrow$ 3.7 $\downarrow$ 0.3
LEO-Mini-Llama-8B 64 1583.0 77.0 75.8 64.5 69.3 75.1 84.5 90.3 38.8
$\Delta$uparrow 106.2 $\uparrow$ 9.5 $\uparrow$ 15.6 $\uparrow$ 2.7 $\uparrow$ 10.8 $\uparrow$ 16.0 $\uparrow$ 12.5 $\uparrow$ 4.0 $\uparrow$ 4.4
LLM
Model
# V tokens
MMEP
MMBench
SEEDI
GQA
SQA
MMMU
POPE
AI2D
VizWiz
TextVQA
DocVQA
ChartQA
OCRBench
Vicuna-7B [6] & Qwen-7B [2] InstructBLIP [8] 32 - - - 49.2 60.5 - - - 34.5 50.1 - - -
LLaVA-1.5 [39] 576 1511 64.3 66.1 62.0 66.8 - 85.9 - 50.0 58.2 - -
LLaVA-NeXt [40] 2880 1519 - - 64.2 70.1 35.1 - 66.6 57.6 64.9 74.4 54.8 -
InternVL [5] 1792 1525 64.3 - 62.9 - - 86.4 - 52.5 57.0 - - -
VILA [37] 576 1533 68.9 61.1 62.3 68.2 - 85.5 - 57.8 64.4 - - -
Monkey [35] 256 - - - 60.7 69.4 - - 62.6 61.2 67.6 66.5 65.1 -
MoVE-based MLLMs
Brave-X5 [25] 160 - - - 52.7 - 87.6 - 54.2 - - - -
Mini-Gemini [34] 576 1523 65.8 - 64.5 71.1 36.1 - - - 65.2 - - -
Mousi-X3 [11] 576 - 66.8 66.0 63.3 70.2 - 87.3 - - 58.0 - - -
LEO [1] 512 - 72.9 72.2 64.8 78.5 36.4 88.0 69.6 57.9 68.8 80.1 71.0 -
DeepSeek-VL [45] 576 - 73.2 70.4 - - 36.6 88.1 - - - - - 45.6
Eagle-X5 [54] 1024 1528 68.4 73.9 64.9 69.8 36.3 88.8 - 54.4 71.2 78.6 67.8 52.9
Llama-8B [44] Cambrian-1 [57] 576 1547 75.9 74.7 64.6 80.4 42.7 - 73.0 - 71.7 77.8 73.3 62.4
LLaVA-NeXT [40] 2880 1604 72.5 72.7 65.2 72.8 41.7 - 71.6 - 64.6 72.8 69.5 49.0
MoVE-based MLLMs
Mini-Gemini-HD [34] 2880 1606 72.7 73.2 64.5 75.1 37.3 - 73.5 - 70.2 74.6 59.1 47.7
Eagle-X4-Plus [54] 1024 1559 75.9 76.3 64.9 84.3 43.4 - 76.1 - 77.1 86.6 80.1 62.6
4 Experiments
4.1 Implementations and Benchmarks
Benchmarks. We evaluate LEO-Mini on several image understanding tasks [14, 41, 29, 22, 46, 64, 32, 26, 15, 16, 55, 49, 48]. Details are deferred to the Appendix.
Baselines. We compare LEO-Mini with several baseline methods, including general MLLMs [8, 39, 40, 5, 37, 35, 47, 34, 11, 1, 45, 57, 54] and token reduction MLLMs [20, 52, 33, 63, 31, 68].
Models. We use Vicuna-v1.5-7B [6] and Llama3-8B [44] as the LLM. For vision experts, we follow the general design of EAGLE [54] and use CLIP [50], ConvNeXt [43], Pix2Struct [28] and EVA-02 [12] for LEO-Mini-Llama-8B. Similarly, we add another vision expert SAM [27] for LEO-Mini-Vicuna-7B. The visual projector is a 2-layer MLP with the GELU activation function [18]. MMoE-LLM is only applied to the MLP in each block of the LLM. We use 3 experts for MMoE-LLM with $k=1$ and $1$ general expert being consistently activated. Each expert is a LoRA block [19] with rank of $16$. The routing network is a 2-layer MLP with the GELU activation function.

### v2

@@ 3.3 Mixture of Multi-modal Experts (MMoE)
4 Experiments
4.1 Implementations and Benchmarks
Models. We use Vicuna-v1.5-7B Chiang et al. (2023) and Llama3-8B Llama Team (2024) as the LLM. For vision experts, we follow the general design of EAGLE Shi et al. (2024) and use CLIP Radford et al. (2021), ConvNeXt Liu et al. (2022), Pix2Struct Lee et al. (2023) and EVA-02 Fang et al. (2024c) for LEO-Mini-Llama-8B. Similarly, we add another vision expert SAM Kirillov et al. (2023) for LEO-Mini-Vicuna-7B. The visual projector is a 2-layer MLP with the GELU activation function Hendrycks and Gimpel (2023). MMoE-LLM is only applied to the MLP in each block of the LLM. We use 3 experts for MMoE-LLM with $k=1$ and $1$ general expert being consistently activated. Each expert is a LoRA block Hu et al. (2022) with rank of $16$. The routing network is a 2-layer MLP with the GELU activation function.
More details on training and evaluation are deferred to the Appendix.
4.2 Main Results
We compare our LEO-Mini with several state-of-the-art token reduction MLLMs (table˜1). LEO-Mini outperforms all the token reduction MLLMs on all tasks. Specifically, LEO-Mini with Llama3-8B achieves 1583.0 on MME (perception), 77.0 on MMBench, 75.8 on SeedBench, 64.5 on GQA, 69.3 on VizWiz, 75.1 on TextVQA, 84.5 on SQA, 90.3 on POPE, and 38.8 on MMMU, outperforming the best baseline by a large margin. This demonstrates the effectiveness of LEO-Mini in improving the performance of MLLMs on various tasks. Moreover, LEO-Mini with Vicuna-7B also outperforms the best baseline on all tasks except for GQA and MMMU, showing the generalization ability of LEO-Mini with different LLMs.
Figure 5: Visualization of the expert choice using LEO-Mini-Llama-8B on TextVQA, ScienceQA, and GQA. Best viewed in color. “E” refers to Experts.
4.3 Ablation Studies
In this section, we conduct ablation studies to analyze the effectiveness of different components in LEO-Mini, the number of visual tokens, and the amount of SFT data using LEO-Mini-Llama-8B. Ablations using other token reduction modules and MoE are presented in the Appendix.
Is 1 visual token enough for MLLMs? To understand whether 1 visual token is enough for representing the visual information, we conduct the experiments and present the results in Table˜2. LEO-Mini with 1 visual token ($N^{V}=1$) shows a slight decrease across some metrics compared to LEO-Mini with 64 tokens, which is reasonable. For example, in the MMEP, the score dropped from 1583 to 1565, and in SEEDI, the performance slightly decreased from 75.8 to 75.6. Similarly, in the domains of GQA and SQA, the scores decrease from 64.5 to 64.2 and 84.5 to 83.6, respectively. This trend continues across other evaluated areas such as MMMU, POPE, AI2D, TextVQA, ChartQA, and OCRBench. While the use of only 1 visual token consistently leads to lower performance, it greatly improves model’s efficiency as shown in Table˜4. To further understand how increasing the number of visual tokens impacts the performance, we conduct experiments with 16 and 256 visual tokens. The results show that the performance on most of the benchmarks is improved as the number of visual tokens increases. This is reasonable as more visual tokens can provide more detailed visual information to the model, which can help the model to better understand the visual information. However, we also observe that the performance on some benchmarks is slightly decreased, e.g., AI2D, which might be due to the overfitting issue.
LLM Model # V tokens MMEP MMBench SEEDI GQA SQA MMMU POPE AI2D VizWiz TextVQA DocVQA ChartQA OCRBench
Vicuna-7B& Qwen-7B
InstructBLIP Dai et al. (2023) 32 - - - 49.2 60.5 - - - 34.5 50.1 - - -
LLaVA-1.5 Liu et al. (2024b) 576 1511 64.3 66.1 62.0 66.8 - 85.9 - 50.0 58.2 - -
LLaVA-NeXt Liu et al. (2024c) 2880 1519 - - 64.2 70.1 35.1 - 66.6 57.6 64.9 74.4 54.8 -
InternVL Chen et al. (2024c) 1792 1525 64.3 - 62.9 - - 86.4 - 52.5 57.0 - - -
VILA Lin et al. (2024b) 576 1533 68.9 61.1 62.3 68.2 - 85.5 - 57.8 64.4 - - -
Monkey Li et al. (2024f) 256 - - - 60.7 69.4 - - 62.6 61.2 67.6 66.5 65.1 -
MoVE-based MLLMs
LLaVA-HR Luo et al. (2024) 1024 1554 - 64.2 64.2 65.1 - 87.6 - 48.7 67.1 - -
Brave-X5 Kar et al. (2024) 160 - - - 52.7 - 87.6 - 54.2 - - - -
Mini-Gemini Li et al. (2024e) 576 1523 65.8 - 64.5 71.1 36.1 - - - 65.2 - - -
Mousi-X3 Fan et al. (2024) 576 - 66.8 66.0 63.3 70.2 - 87.3 - - 58.0 - - -
LEO Azadani et al. (2025) 512 - 72.9 72.2 64.8 78.5 36.4 88.0 69.6 57.9 68.8 80.1 71.0 -
DeepSeek-VL Lu et al. (2024) 576 - 73.2 70.4 - - 36.6 88.1 - - - - - 45.6
Eagle-X5 Shi et al. (2024) 1024 1528 68.4 73.9 64.9 69.8 36.3 88.8 - 54.4 71.2 78.6 67.8 52.9
Llama-8B
Cambrian-1 Tong et al. (2024) 576 1547 75.9 74.7 64.6 80.4 42.7 - 73.0 - 71.7 77.8 73.3 62.4
LLaVA-NeXT Liu et al. (2024c) 2880 1604 72.5 72.7 65.2 72.8 41.7 - 71.6 - 64.6 72.8 69.5 49.0
MoVE-based MLLMs
Mini-Gemini-HD Li et al. (2024e) 2880 1606 72.7 73.2 64.5 75.1 37.3 - 73.5 - 70.2 74.6 59.1 47.7
Eagle-X4-Plus Shi et al. (2024) 1024 1559 75.9 76.3 64.9 84.3 43.4 - 76.1 - 77.1 86.6 80.1 62.6
[…]
Figure 6: Qualitative results across four vision-language tasks demonstrating LEO-MINI’s detailed visual understanding. The images are taken from TextVQA Singh et al. (2019), DocVQA Mathew et al. (2021), and MMMU Yue et al. (2024).
How does MMoE-LLM switch between different LORA experts? We visualize the expert choice in fig.˜5 with GQA (General), TextVQA (OCR), and SQA (Knowledge). We observe that the model effectively switches between different LoRA experts based on the input data. For example, in the SQA task, the model mainly activates expert 2 at layer 0, while for TextVQA and GQA, the model evenly activates three experts. When it comes to the final layer, such as layer 29, the model evenly activates experts 0 and 1 for TextVQA and GQA, while for SQA, the model mainly activates expert 1, respectively. This indicates that MMoE-LLM effectively utilizes the multi-modal input instructions to switch between different experts. We also compare the expert choice between LEO-Mini and the vanilla MoE. Results are presented in Section˜A.3.
4.4 How Does LEO-Mini Compare to General MLLMs?
We also compare LEO-Mini to general and MoVE-based MLLMs (Table˜3). With Vicuna-7B, LEO-Mini outperforms all the general and MoVE-based MLLMs on MME, POPE, AI2D, and OCRBench by 15, 1.2, 2.6, and 2.7 points, respectively, with only 64 visual tokens. LEO-Mini with Vicuna-7B also achieves the second best performance on SeedBenchI, SQA, and TextVQA. On the other side, with Llama3-8B, LEO-Mini achieves stronger performance as it outperforms all the general and MoVE-based MLLMs on MMBench, SQA, and CharQA by 1.1, 0.2, and 0.4 points, respectively, while reducing the number of visual tokens by 97.77% compared to LLaVA-NeXt and Mini-Gemeni-HD. Moreover, compared to the general MLLMs, LEO-Mini with Llama3-8B achieves comparable performance on other benchmark datasets as the discrepancy is marginal.
4.5 Qualitative Analysis
To understand the detailed visual understanding of LEO-Mini, we conduct a case study for qualitative analysis on four vision-language tasks Singh et al. (2019); Mathew et al. (2021); Yue et al. (2024) as shown in Figure˜6. We use LEO-Mini-Llama3-8B with 64 tokens.
4.6 Efficiency Analysis
To understand how efficient LEO-Mini is, we compare the number of visual tokens, FLOPs, and CUDA processing time of LEO-Mini with other We also compare LEO-Mini to MLLMs Liu et al. (2024c); Shi et al. (2024); Tong et al. (2024) using Llama3-8B. Specifically, we use Pytorch Profiler to measure the FLOPs, CUDA time, and GPU memory usage using the figure and “What is shown in this image?” as input. We use one A6000 GPU with 48 GB memory to inference the models.
The comparison is shown in Table˜4. LEO-Mini with Llama3-8B is 97.77% more efficient in terms of visual tokens, 63.83% in terms of FLOPs, and 90.69% in terms of CUDA time compared to LLaVA-Next. Moreover, as shown in Figure˜1, LEO-Mini outperforms LLaVA-Next by 6% on average performance with only 64 visual tokens. Even compared to the most powerful MLLM with similar performance, i.e., Eagle-X4-Plus, LEO-Mini is 93.75% more efficient in terms of visual tokens, 44.10% for FLOPs, and 68.40% for CUDA time. More excitingly and importantly, LEO-Mini with only 1 token achieves comparable performance to the state-of-the-art MLLMs with thousands of visual tokens with faster inference time and lower computational cost in terms of FLOPs.
Models # of VT FLOPs (T) CUDA Time (s) GPU Memory (GB)
LLaVA-Next Liu et al. (2024c) 2880 45.5 7.037 14.74
Eagle-X4-Plus Shi et al. (2024) 1024 29.4 2.073 20.10
LEO-Mini-Llama-8B 1 13.7 0.593 19.28
LEO-Mini-Llama-8B 64 16.4 0.655 19.34
$\Delta$ $\downarrow$ 97.77% $\downarrow$ 63.83% $\downarrow$ 90.69%
5 Conclusion
In this paper, to address the redundancy of visual tokens in MLLMs, we propose a novel MLLM, LEO-Mini, that significantly reduces the number of visual tokens while boosting visual reasoning capabilities. LEO-Mini incorporates a novel token reduction module, CoTR, to consolidate a large number of visual tokens into a smaller set of tokens, using the similarity between visual tokens, text tokens, and a compact learnable query. However, simply reducing the number of visual tokens leads to an information loss. To avoid the loss and boost the visual comprehension ability with minimal computational overhead, LEO-Mini employs a novel mixture of multi-modal experts module, MMoE, that includes a set of language (LoRA) experts and a set of vision experts trained on diverse domain-specific data. For better switching between different LoRA experts, MMoE employs a new router that takes the text and visual tokens as additional inputs. MMoE also includes a general LoRA expert that is always activated to learn general knowledge. We evaluate LEO-Mini on various vision-and-language tasks, showcasing its potential for practical applications with improved efficiency and performance compared.
6 Limitations
First, the proposed CoTR needs training. Introducing a training-free token reduction module might be a promising direction. Second, the proposed MMoE is designed to be efficient and scalable, but it may not be optimal for all tasks. For vision experts, due to the computational limitations, we use a fixed set of experts. It would be interesting to explore more vision experts from a wide range of domains and introduce more advanced expert selection mechanisms, such as dynamic vision expert routing. Last, due to the computational limitations, we did not test the efficiency of LEO-Mini using bigger LLMs with 13B or 67B parameters.
7 Potential Risks
LEO-Mini is an MLLM that can be used for various tasks, including visual question answering, image captioning, and document understanding. However, as an inherent risk of (multi-modal) LLMs, LEO-Mini may generate biased or harmful content, especially when the query data contains sensitive information. We recommend that users carefully review the generated content and apply appropriate filters to mitigate potential risks.
Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC)-CSE Research Community project entitled “An End-to-End Approach to Safe and Secure AI Systems” and NSERC’s Postdoctoral Fellowship. Researchers funded through the NSERC-CSE Research Communities Grants do not represent the Communications Security Establishment Canada or the Government of Canada. Any research, opinions, or positions they produce as part of this initiative do not represent the official views of the Government of Canada.
Appendix A Experiments
A.1 Benchmark Datasets
We evaluate our method on the following benchmark datasets: MME Fu et al. (2024), MMBench Liu et al. (2024d), Seed-Bench Li et al. (2024a), GQA Hudson and Manning (2019), SQA Lu et al. (2022), MMMU Yue et al. (2024), POPE Li et al. (2023), AI2D Kembhavi et al. (2016), VizWiz Gurari et al. (2018), TextVQA Singh et al. (2019), DocVQA Mathew et al. (2021), ChartQA Masry et al. (2022), and OCRBench Liu et al. (2024e).
MME Fu et al. (2024). The MME benchmark is designed to rigorously evaluate a model’s perceptual and cognitive abilities through 14 subtasks. It employs carefully constructed instruction-answer pairs and concise instructions to minimize data leakage and ensure fair evaluation. This setup provides a robust measure of a model’s performance across various tasks.
MMBench Liu et al. (2024d). MMBench offers a hierarchical evaluation framework, categorizing model capabilities into three levels. The first level (L-1) focuses on perception and reasoning. The second level (L-2) expands this to six sub-abilities, while the third level (L-3) further refines these into 20 specific dimensions. This structured approach allows for a nuanced and comprehensive assessment of a model’s multifaceted abilities.
Seed-Bench Li et al. (2024a). SEED-Bench consists of 19K multiple-choice questions with accurate human annotations, covering 12 evaluation dimensions including both the spatial and temporal understanding.
GQA Hudson and Manning (2019). GQA is structured around three core components: scene graphs, questions, and images. It includes not only the images themselves but also detailed spatial features and object-level attributes. The questions are crafted to assess a model’s ability to comprehend visual scenes and perform reasoning tasks based on the image content.
ScienceQA Lu et al. (2022). ScienceQA spans a wide array of domains, including natural, language, and social sciences. Questions are hierarchically categorized into 26 topics, 127 categories, and 379 skills, providing a diverse and comprehensive testbed for evaluating multimodal understanding, multi-step reasoning, and interpretability.
@@ A.1 Benchmark Datasets
MMMU Yue et al. (2024). MMMU includes 11.5K meticulously collected multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. These questions span 30 subjects and 183 subfields, comprising 30 highly heterogeneous image types, such as charts, diagrams, maps, tables, music sheets, and chemical structures.
POPE Li et al. (2023). POPE is tailored to assess object hallucination in models. It presents a series of binary questions about the presence of objects in images, using accuracy, recall, precision, and F1 score as metrics. This approach offers a precise evaluation of hallucination levels under different sampling strategies.
AI2D Kembhavi et al. (2016). AI2D is a dataset of over 5000 grade school science diagrams with over 150000 rich annotations, their ground truth syntactic parses, and more than 15000 corresponding multiple choice questions.
VizWiz Gurari et al. (2018). VizWiz consists of over 31,000 visual questions originating from blind people who each took a picture using a mobile phone and recorded a spoken question about it, together with 10 crowdsourced answers per visual question.
TextVQA Singh et al. (2019). TextVQA emphasizes the integration of textual information within images. It evaluates a model’s proficiency in reading and reasoning about text embedded in visual content, requiring both visual and textual comprehension to answer questions accurately.
DocVQA Mathew et al. (2021). DocVQA consists of 50,000 questions defined on 12,000+ document images.
ChartQA Masry et al. (2022). CharQA is a large benchmark covering 9.6K human-written questions as well as 23.1K questions generated from human-written chart summaries.
OCRBench Liu et al. (2024e). OCRBench is a comprehensive benchmark for evaluating the OCR capabilities of multi-modal language models across five key tasks: text recognition, scene text-centric and document-oriented VQA, key information extraction, and handwritten mathematical expression recognition.
A.2 Training Details
The training of LEO-Mini consists of three stages, as shown in Table˜6. Stage 1: Warming up for the visual projector. This stage pre-trains the visual projector while keeping the LLM and vision experts frozen. The LLM and vision experts are initialized from the base models, while the vision projector is randomly initialized. Stage 2: Supervised fine-tuning. In this stage, we fine-tune the entire model, including the LLM, vision experts, and visual projector. Stage 3: Supervised fine-tuning for reducing the number of visual tokens. In this stage, we introduce CoTR to the model and perform LoRA fine-tuning (MMoE-LLM) for efficiency. Specifically, CoTR and MMoE-LLM are randomly initialized and fine-tuned, while the LLM and vision experts remain frozen. To avoid the risk of routing collapse Shazeer et al. (2017), we employ a balanced loss for MMoE with a hyperparameter $\lambda=0.05$, following previous works Shazeer et al. (2017); Tian et al. (2024).